OpenStudy (falling_in_katt):

***WILL MEDAL*** Perform the operation indicated, then place the answer in the proper location on the grid. Write your answer in descending powers of m. (5m ^ 3 + 3mn ^ 2) + (-7m ^ 2 n + mn ^ 2 - n ^ 3) - (mn ^ 2 + 2m ^ 3)

OpenStudy (amoodarya):

\[(5m ^ 3 + 3mn ^ 2) + (-7m ^ 2 n + mn ^ 2 - n ^ 3) - (mn ^ 2 + 2m ^ 3)=\\(5-2)m^3-7m^2n+mn^2(3+1-1)-n^3\\3m^3-7m^2n+3mn^2-n^3\]

OpenStudy (swamijatin2):

(5m^3 + 3mn^2) + (-7m^2n + mn^2 - n^3) - (mn^2 + 2m^3) = (5 + 2)m^3 + (3 + 1 - 1)mn^2 + (-7)m^2n + (-1)n^3 = 7m^3 + 3mn^2 - 7m^2n - n^3